OpenStudy (mathmath333):

the units digit of a two digit number is greater than its ten digit by 2,and the product of that number by the sum of its digit is 144. find the number.

OpenStudy (mathmath333):

a.14 b.24 c.46 d.35 e.20

OpenStudy (anonymous):

ok so for the first part you only have to options correct?

OpenStudy (anonymous):

your options are b and d so now you have to figure out which is the final answer

OpenStudy (anonymous):

so according to the question you need to add the ten unit and the one unit so 2+4=6 4+6=10 then using those sums find the product between the sums and the original number

OpenStudy (anonymous):

since the answer for that area is given at 144... 46*10 would not make sense that leaves the answer to be 24*6 (=144) so your answer is B.